OpenStudy (anonymous):

What are the real zeros, because 1 -2 -3 don't seem to be working!? I'm trying to find the real zeros to this problem but I just ******* cant! f(x) = x^3 + 4x^2 + x − 6

OpenStudy (anonymous):

Those are the real zeros...

OpenStudy (anonymous):

\[f(1)=1^3+4(1)^2+1-6=0\] \[f(-2)=(-2)^3+4(-2)^2+(-2)-6=0\] \[f(-3)=(-3)^3+4(-3)^2+(-3)-6=0\]

OpenStudy (anonymous):

So it's a problem on the website then